OpenStudy (anonymous):

Please hjäp me to find the second derivative f''(-1) to f (x) = In (x ^ 2 +2 x +6)

OpenStudy (anonymous):

f'(x) = 2(x+1)/(x^2+2x+6) f''(x)=-2(x^2+2x-4)/(x^2+2x+6)^2 Therefore f''(-1)=2/5

myininaya (myininaya):

f'(x)=(2x+2)/(x^2+2x+6) now use the quotient rule to find second derivative f''(x)={2(x^2+2x+6)-(2x+2)(2x+2)}/{(x^2+2x+6)^2} ={2x^2+4x+12-(4x^2+8x+4)}/{(x^2+2x+6)^2} ={-2x^2-4x+8}/{(x^2+2x+6)^2}

OpenStudy (anonymous):

\[f(x) = \ln(x^{2} +2x+6) \Rightarrow f'(x)=\frac{1}{x^{2}+2x+6}(2x+2) = \frac{2x+2}{x^{2}+2x+6}\] Now quotient rule: \[f''(x) = \frac{(2)(x^{2}+2x+6)-(2x+2)(2x+2)}{(x^{2}+2x+6)^{2}}\]

OpenStudy (anonymous):

Thanks, thanks and thanks